### v3

@@ 2.2.6 AI-based Situational Awareness and Fault Detection
Network situational awareness technology is capable of endowing mobile communication networks with real-time monitoring, prediction, and response capabilities. Its application range spans various aspects, including network management and optimization, security monitoring, as well as fault detection and prediction [104]. The AI-based situational awareness system can ensure the security of all links within the 6G network. It can prevent malicious penetration attacks, data loss, and malicious tampering, guarantee the credibility of access terminals, the reliability of transmission channels, and the controllability of service applications, and achieve panoramic security monitoring, thereby comprehensively enhancing the comprehensive security defense capabilities of all 6G services.
Specifically, the situational awareness technology relying on AI is a data-driven network cognitive paradigm. It is capable of collecting, storing, and analyzing vast amounts of network data via big data platforms and DL techniques, thereby providing precise threat intelligence and full traffic inspection [105]. This technology employs AI approaches such as DL to precisely identify the states of various elements including network traffic, logs, and network KPI, predict network development trends, and formulate accurate response strategies through the analysis of perceived information and predicted states [106], realizing the shift from “passive defense” to “active defense”. The AI-driven threat representation and causal reasoning can automatically identify and infer the origins of attacks, preventing the recurrence of similar security events. The real-time threat alert and response handling capabilities assist users in quickly grasping attack circumstances, affected devices, and emergency handling tactics, thus ensuring the visualization and continuous monitoring of network security postures.
The network fault analysis constitutes one of the crucial elements in the O&M of mobile communication networks. As the number of mobile communication access devices increases and the network scale expands, the quantity of network alarms is growing explosively, and the relationships among them are highly intricate [107]. Employing AI to build a fault detection analysis model endows the model with the capacity for high-dimensional data analysis, enabling it to serve as a more intelligent and efficient tool for fault analysis [108]. The fault detection method based on AI automatically extracts key features from massive network alarm data through DL and ML algorithms. It combines network topology structures, configuration settings, KPI, and historical alarm and fault handling data to construct a fault analysis model. Through conducting a fusion analysis of these data, the system can effectively identify common issues within the network and perform accurate fault diagnoses. To enhance the system’s adaptability and precision, the results obtained after fault handling will be utilized as feedback to update and optimize the fault diagnosis model, allowing the system to operate efficiently under constantly changing network environments and business requirements [109]. The specific operational workflow is illustrated in Figure 12.
@@
By integrating AI technology, we can establish a systematic and intelligent network fault analysis and tracing back scheme, which can provide automated, accurate and flexible fault detection and tracing back capabilities and continuously enhance the efficiency of fault diagnosis and handling. Meanwhile, it can also adapt to the changes of network environment and business requirements.
@@ 3 Network for AI
At present, preliminary achievements have been made in promoting AI4NET. A series of AI use cases and scenarios have been developed, and part of the standardization work has been accomplished under the leadership of 3GPP. This has led to the formation of various module-level “plug-in” AI functions in 5G networks, which are used to optimize the efficiency of wireless transmission and improve system performance for identified and specific communication problems. Although this approach is convenient, it still lacks considerations of systematicity and interpretability, which greatly limits the scalability of AI capability and the ability of model generalization, and hinders the development of more efficient AI services in the networks. Essentially, this is because the current network architecture cannot provide sufficient support for the in-depth integration of AI into the network.
Although key technologies such as cloud-native, information technology (IT) software-defined, and service virtualization have been developed during the evolution of 5G networks (for example, CN that supports service-based architecture and cloud-based deployment of CN functions, etc.), and 3GPP has standardized and defined the Network Data Analytics Function (NWDAF) as AI functional components in the CN, along with relevant mechanism processes, and is attempting to introduce and define standardized AI operation paradigms. However, these efforts are still restricted by multiple conditions including technology maturity, security, and the complexity of system O&M. As a result, the current access network architecture generally still retains the traditional “siloed communication-technicalized BS architecture”, i.e., the current BS architecture is still highly integrated and designed based on communication services, and cannot effectively connect and adapt to emerging AI services. This leads to the imperfection of the overall network design, with relatively major loopholes and deficiencies remaining. Because from the perspective of the end-to-end AI service process, there is no corresponding definition of AI functional components on the access network side, and the docking and integration of AI service functions between CN and the access network have not been achieved, resulting in the current network architecture being unable to support the intelligent vision of 6G.
To support “native intelligence” and achieve ubiquitous general intelligence, the 6G network requires groundbreaking innovative innovations in the network architecture to enable the endogenous embedding of AI capabilities within the network. This represents the core concept of Phase II (i.e., NET4AI) in the deep integration of 6G and AI, signifying a shift from a communication network solely providing connectivity services to a new network paradigm that offers a diverse array of services, including connectivity, computing power, data, and algorithms. This new paradigm would span CNs, access networks, and terminals, to provide comprehensive support of AI. This enables AI training/inference to be executed more efficiently in real time while enhancing data security and privacy protection, among other benefits.
@@ 3.1 NET4AI Architecture over 6G
The research on architecturally integrating 6G networks with AI has commenced and is under continuous exploration and development. Both industry and academia have shown immense interest, with preliminary visions of novel concepts and target architectures tailored for the proposed fusion of 6G and AI.
@@
China Mobile in [110] proposed a systematic 6G network architecture design that transcends connectivity in the industry, centered on the core concept of “three layers and six planes”. The “three layers” refers to the physical resource layer, the network function layer, and the application & service layer, while the “six planes” comprise the communication plane, the data plane, the computing plane, the intelligence plane, the security plane, and the management&orchestration plane. The first five planes are independent functional planes formed by the decoupling of the 6G network from end to end, providing corresponding capabilities and services. The management&orchestration plane realizes the flexible combination of functions across layers and planes, realizes the management and scheduling of multiple resources, and forms a completes service chain. This architecture focuses on multi-dimensional capability elements within the network, aiming to provide comprehensive services throughout the information flow process. In [111], China Telecom proposed a ubiquitous and ultra-converge 6G network architecture, built upon the core concept of “three layers and three sectors”. The “three layers” consist of the Infrastructure Layer, the Network Function Layer, and the Network Operating System Layer, while the “three sectors” embody Data Integration, Intelligent Simplicity, and Trustworthiness. This architecture leverages the Intelligent Simplicity sector to integrate the intelligent brain, connections, and facilities across the three layers, thereby enabling intelligent inclusiveness. China Unicom in [112] envisioned furture 6G network as intelligent, converged, green, and trustworthy. Its architecture is structured vertically into three communication layers of land-based, air-based, and space-based and horizontally into three domains of Perception Resource Domain, Function Control Domain, and Service Application Domain. These layers and domains are interconnected by two chains: Intelligent Native and Secure & Trustworthy.
@@
Huawei in [113] unveiled a task-centric AI architecture of 6G networks, asserting that 6G networks will introduce new resource dimensions and support the coordination of multi-dimensional resources. This design allows 6G networks to natively support AI and achieve a transformation from session-centric to task-centric operations. CICT in [114, 115] put forward the concept of multi-network integration and ubiquitous intelligence, envisioning that 6G will realize global three-dimensional deep coverage, develop body area networks centered on human beings, form multi-layer coverage with both breadth and depth and learn the ubiquitous intelligence system with multi-network integration. OPPO in [116] suggested that 6G will redefine human interaction and AI, by making AI a ubiquitous infrastructure. The 6G system architecture will emphasize the coordination across intelligence, performance, and flexibility, and integrate deeply AI capabilities into the user plane, control plane, and function plane. More authoritative and official, The IMT-2030 (6G) Promotion Group in China envisioned the 6G network architecture in [117], as an open and innovative platform for information services, offering capabilities that transcend mere connectivity. These capabilities encompass computing power networks, trusted security, sensing, data services, and AI-based network intelligent autonomy.
@@
Ericsson in [118] is focused on researching cognitive networks and leveraging AI to achieve data-driven security and highly automated network operations. Qualcomm advocated that AI and ML will fundamentally transform the design and deployment of wireless communication and networking systems. As the leader of the EU’s 6G project Hexa-X, Nokia in [119] stated that AI will be the primary driving force behind the technological transformation of the NR interface for 6G. The company is creating unique AI use cases and scenarios and developing foundational AI technologies. LG Electronics is focused on the research and development of 6G AI technology, anticipating that 6G systems will usher in the era of the “Internet of Everything and Environment” powered by AI. Japan’s “Beyond 5G Promotion Strategy” project aimed to introduce AI into network management, enabling network autonomy and driving network evolution towards intelligence.
@@
Following the industrial perspectives on AI integrated 6G network architecture, academic research has also proposed various innovative ideas for shaping the future 6G network architecture. The authors put forth a 6G network design philosophy of “Human-Machine-Thing-Spirit”[120], emphasizing the inevitable trend of AI integration in 6G, leveraging cognitive enhancement and decision-making simulation to intelligently define network requirements, ensuring secure and reliable network transmission, and realizing intent-driven networks. The authors of [121] introduced a fusion of 6G RAN and AI, identifying four key characteristics as green, multi-dimensional, stereo, and full-scenario service. The authors proposed that 6G will integrate communication, computation, control/caching, and AI to support full-application scenarios, and will be inherently secure[122]. The authors of [123] demonstrated a vision of integrating 6G and AI and studied implementation examples in wireless communications based on two classic AI algorithms: DL and RL. In [30], the authors argued that AI will become the cornerstone of 6G with “Intelligent inclusion” serving as an essential feature of 6G, enabling the network to provide intelligent services to all types of end users. The authors of [124] advocated that AI will be used in wireless communications and change the top-level architecture of wireless networks. The authors of [125] emphasized that the NET4AI paradigm should leverage the capabilities of 6G networks, such as precise wireless sensing, high data-rate transmission, and ubiquitous connectivity, for intelligence distillation at the network edge, promoting edge AI and, therefore enabling ubiquitous and trustworthy AI applications in wireless networks. Among them, federated edge learning (FEEL) is expected to become an important learning architecture over 6G. In [126, 127], the authors pointed out that in the 6G era, networks and AI will be fully integrated, and customized AI services will be provided on the wireless network side. The authors of [128] reported that 6G would enhance its progress in intelligence, especially in edge intelligence. Among them, integrating DT and edge networks will be a promising attempt. The authors of [129] initiated the 6G Flagship research program, advocating for realizing context-aware intelligent services and applications for human and non-human users through AI.
@@
Incorporating the aforementioned research ideas and architectural visions, the industry has reached a core understanding of the intelligent design of 6G network architectures, emphasizing “distributed structure”, “endogenous intelligence”, and “integrated simplicity”. Based on this consensus, future networks will evolve into advanced, integrated platforms that surpass mere communication connectivity services, offering multi-dimensional services and corresponding network capabilities beyond communication. Generally speaking, the consensus suggests that 6G networks are expected to provide communication connectivity, data capability, computing capability, security capability, AI capability, management and orchestration capability and their corrsponding services. Furthermore, they will cater to diverse businesses by offering end-to-end lifecycle management services throughout their entire process. These consensuses are illustrated in Figure 13.
Figure 13: Logical consensus of 6G network. Based on current research, we can conclude that the 6G system will inherently possess six network capabilities encompassing communication connectivity, data, computation, security, AI, and management & orchestration. We will offer six fundamental service categories that evolve from these capabilities. In the 6G era, any business or scenario can be disaggregated into demands and combinations of these basic services, with the network capable of synergistically and dynamically scheduling multi-dimensional capabilities on demand, thereby supporting all scenarios and businesses. For instance, the autonomous driving business can be decomposed into a combination of six basic services: communication connectivity (data and model transmission between vehicles and vehicles/BSs/servers), data (storage and scheduling of vehicle driving data), computation (data preprocessing and model gradient optimization), security (transmission security, data security, and model security), AI (model training and model aggregation), and management & orchestration (cooperative transmission/computation/training). These services are realized through the corresponding six capabilities provided by the 6G network.
Communication Connectivity Capability: In 6G, the control plane and user plane within the communication layer will undergo further enhancement designs to deepen the service orientation and openness of network functions to address the complex and diverse business scenarios and emerging demands of the future.
(1) Enhanced control plane: Within the control plane, 6G necessitates enhancing and optimizing the service management and interfaces of existing control plane functions to reduce the complexity of network function and signaling interaction processes. Additionally, 6G will extend the conventional control plane by integrating space-terrestrial ubiquitous access control management, enabling new functionalities and service-oriented offerings for session management and policy control across communication, computing, data, and other essential elements.
(2) Enhanced user plane: The enhancement of the user plane focuses on augmenting or introducing new, more generalized, and flexible user plane forwarding protocols that can perform agile service forwarding and processing based on communication, computing, and other elements. Leveraging microservice technologies and next-generation hardware support, e.g., high-performance programmable chips, cloud-native technologies, and the P4 (programming protocol-independent packet processors) advanced language, user plane functions (UPF) will be restructured into service-oriented modules. This flexible invocation and efficient collaboration among functional modules will break the constraints of conventional layered protocols, enhancing data forwarding efficiency and bandwidth utilization.
Data Capability: The conventional user plane architecture struggles to support distributed management and efficient utilization in response to the diverse and highly time-sensitive requirements of sensory data, AI data, network behavior, status data, and other forms. Consequently, 6G networks necessitate a novel data capability that differs from conventional user plane. This capability must adhere to data regulation and supervision requirements while enhancing data analysis and processing efficiency, addressing data management challenges, and enabling value realization.
The data capability supports the realization of information collection, cleaning, storage, and more, enabling the complete lifecycle management of data service processes while ensuring data privacy and security. By leveraging a hybrid data processing and storage model that combines centralized and distributed approaches, the 6G network can flexibly address diverse scenario requirements: 1) For sensitive or high-privacy-level data, local processing and storage can be performed at the RAN/TN/CN and terminal sides to ensure no data leakage; 2) Global and comprehensively analyzed data, on the other hand, is processed and stored centrally in the data center, which also assumes the responsibilities of global data orchestration and access authentication, achieving efficient data organization and collaboration. Through tight cooperation between the central data center and edge nodes, 6G can efficiently manage massive amounts of data originating from various technological and business domains while safeguarding privacy, thereby facilitating the development of data applications and services.
Computation Capability: The 6G computation capability will encompass functionalities such as computational sensing, control, and execution. Among them, the computational sensing function gathers information on the ubiquitous computing power resources distributed across the spectrum from central clouds to terminals. The computational control function, tailored to computing service demands, provides pervasive computing power control and scheduling strategies. The computational execution function, based on these scheduling strategies, offers services like session management, traffic control, and policy execution. Additionally, computational capability needs to synergize with communication connectivity to optimize the efficiency and overall energy consumption of emerging services like AI. Furthermore, introducing computational capability involves the creation of a unified environment enabling continuous computing service operations across terminals, edges, networks, and clouds. This allows the selection of optimal computing nodes for computational tasks based on latency, bandwidth, computing power, and energy efficiency requirements, seamlessly transitioning between terminals, networks, edge nodes, and clouds.
Security Capability: conventional security measures have struggled to meet the complex and ever-changing security demands of 6G. Currently, the design of security mechanisms stands independently from the network architecture, leaving room for improvement in both security and efficiency. Hence, 6G networks necessitate establishing an inherently trustworthy and self-driving security system at the core of the system. The rapid advancements in cloud computing, big data, and AI technologies offer technical support for constructing a 6G network security system. The 6G security system will be structured around three key components: security capabilities, security controls, and security decision-making. The security capability layer deploys security resources (including physical security resources and pooled security capabilities) on demand, enabling flexible scheduling. The security control layer oversees the scheduling of security resources and orchestrates trust and security collaborations. Meanwhile, the security decision-making layer integrates AI analytics capabilities with DT networks to facilitate intelligent security analysis, simulation, and the integration and delivery of security service capabilities. Based on this architecture, 6G networks will foster a trustworthy and inherently secure system characterized by intelligent proactive defense, ubiquitous intelligent collaboration, and robust privacy protection.
AI Capability: Aligned with the trend of network intelligence evolution, the 6G network architecture will support ubiquitous and inclusive AI capabilities. This necessitates a breakthrough from the external or patch-based intelligence of 5G, transforming intelligence into an inherent network ability. Consequently, the AI capabilities in 6G networks refer to those that cater to the network’s O&M and provide AI services to users or externally. By endogenously constructing AI elements (e.g., computing power, algorithms, data, etc.) and under centralized intelligent orchestration, management, control, and distributed intelligent execution, 6G enables the on-demand customization of AI services. The network side will provide lifecycle management functions for AI models and data, encompassing capabilities such as AI data management (e.g., training data reporting, gradient reporting), AI model management (e.g., model generation, storage, selection, distribution, updating, replacement, activation/deactivation, and performance monitoring), and AI computing power allocation (e.g., for training and inference execution).
Management &\&& Orchestration Capability: The immense logical structure and technological ecosystem complicate network O&M management, necessitating high network autonomy in 6G networks. Consequently, there is an urgent need for management and orchestration capabilities to uniformly and dynamically orchestrate and schedule various functional components based on business requirements, enabling self-optimization, self-evolution, and self-healing of the network. The 6G system will introduce external applications and service demands while exposing internal diversified resources and functions. This requires flexible cross-functional flow and scheduling of multi-dimensional resources, such as computing, communication, and storage within the network, offering customized and personalized application services based on virtualization and microservice technologies. For instance, AI capabilities and analyzed data within the network can be opened to third parties to provide services and necessary support, a crowdsourcing behavior that necessitates the support of management and orchestration functions. Furthermore, through joint analysis and prediction of computing resources and network traffic, containerized gNBs can automatically scale up or down based on predicted network traffic loads, enabling dynamic resource allocation and network energy savings. Therefore, management and orchestration capabilities are closely related to all other capabilities, orchestrating and managing resources and functions efficiently and flexibly across all functional layers of the 6G network, achieving joint optimization.
Within the consensus of the 6G network architecture, communication connectivity is responsible for establishing end-to-end communication links and providing precise and highly reliable connectivity services. The data capability oversees data collection, cleansing, processing, and storage within the end-to-end network, offering data subscription and update services to other planes. The computation capability is tasked with scheduling and optimizing various computational resources within the network to handle computational tasks. The security capability provides inherent security support for the entire network, offering security services to customers while opening up security capabilities to external partners or facilitating crowdsourcing of external security capabilities. The AI capability provides an AI engine that, in conjunction with functional requirements across various domains within the network, offers big data analytics, AI algorithms, model training services, and simulation verification of related solutions. AI capabilities can be centralized or distributed, which can be distributed on network elements and terminals to support real-time AI applications, and can also be concentrated on the cloud to implement complex algorithms with high computing requirements. Additionally, the network’s AI capabilities can be opened to the outside world, while external AI capabilities can be introduced into the network, enabling the crowdsourcing of AI capabilities. The management & orchestration capability is responsible for scheduling various capabilities within the network to achieve flexible organization, fulfilling the orchestration and management of different resources and functions to meet users’ personalized service demands.
@@
Drawing upon the aforementioned consensus on network capabilities in 6G, we further elaborate on the NET4AI architecture, i.e., the support provided by the other five capabilities/services in the network (i.e., communication connectivity, data, computation, security, and management & orchestration) for the AI capabilities/services, with an emphasis on their connectivity with the AI, as illustrated in Figure 14.

### v4

@@ 2.2.6 AI-based Situational Awareness and Fault Detection
Network situational awareness technology endows mobile communication networks with real-time monitoring, prediction, and response capabilities. Its applications include network management and optimization, security monitoring, as well as fault detection and prediction [109]. AI-based situational awareness collects, stores, and analyzes vast amounts of network data via big data platforms and DL techniques, thereby providing precise threat intelligence and full traffic inspection [110]. This technology employs DL to precisely identify the states of various elements including network traffic, logs, and network key performance indicator (KPI), predict network development trends, and formulate accurate response strategies through the analysis of perceived information and predicted states [111], realizing the shift from “passive defense” to “active defense”. The AI-driven threat representation and causal reasoning can identify and infer the origins of attacks, preventing the recurrence of similar security events.
The network fault analysis constitutes one of the crucial elements in the O&M of mobile communication networks. As the number of mobile devices increases and the network scale expands, the quantity of network alarms is growing explosively, and the relationships among them are highly intricate [112]. Employing AI to build a fault detection analysis model endows the model with the capacity for high-dimensional data analysis, enabling it to serve as a more intelligent and efficient tool for fault analysis [113]. By conducting a fusion analysis of these data, the system can effectively identify common issues within the network and perform accurate fault diagnoses. The results obtained after fault handling will be utilized as feedback to update and optimize the fault diagnosis model, allowing the system to operate efficiently under constantly changing network environments and business requirements [114]. The specific operational workflow is illustrated in Figure 11.
@@
By integrating AI technology, we can establish a systematic and intelligent network fault analysis and tracing back scheme, which can provide automated, accurate, and flexible fault detection and tracing back capabilities and continuously enhance the efficiency of fault diagnosis and handling. Meanwhile, it can also adapt to the changes in network environment and business requirements.
@@ 3 Network for AI
A series of AI use cases and scenarios have been developed, and part of the standardization work has been accomplished under the leadership of 3GPP. This has led to the creation of various module-level “plug-in” AI functions in 5G networks. Although this approach is convenient, it still lacks considerations of systematicity and interpretability, limiting the scalability of AI capability and the ability to model generalization.
Key technologies such as cloud-native architectures, software-defined information technology, and service virtualization have advanced significantly in 5G (e.g., CN supporting service-based architecture and cloud-based deployment of CN functions). 3GPP has also standardized the NWDAF as an AI-enabled component within the CN. This includes defining relevant mechanisms and processes while introducing standardized AI operational paradigms. However, these advancements remain constrained by several factors, including technological maturity, security concerns, and the complexity of system O&M. Consequently, the current access network architecture largely retains the traditional “siloed communication-technicalized BS design.” In this architecture, BSs are highly tailored to communication services, lacking the flexibility to interface with emerging AI-driven services. From an end-to-end AI service perspective, there is no clear definition of AI functional components on the access network side. The integration and coordination of AI service functions between the CN and the access network remain unaddressed. This disconnect hinders the current network architecture from supporting the intelligent vision envisioned for 6G.
To support “native intelligence” and achieve ubiquitous general intelligence, the 6G network requires groundbreaking innovations in the network architecture to enable the endogenous embedding of AI capabilities within the network. This represents the core concept of Phase II (i.e., NET4AI) in the deep integration of 6G and AI, signifying a shift from a communication network solely providing connectivity services to a new network paradigm that offers a diverse array of services, including connectivity, computing power, data, and algorithms. This new paradigm would span CN, access networks, and terminals, to provide comprehensive support of AI.
@@ 3.1 NET4AI Architecture over 6G
China Mobile [115] proposed a systematic 6G network architecture design that transcends connectivity in the industry, centered on the core concept of “three layers and six planes”. The “three layers” refer to the physical resource layer, the network function layer, and the application&service layer. The “six planes” comprise the communication plane, the data plane, the computing plane, the intelligence plane, the security plane, and the management&orchestration plane. The first five planes are independent functional planes formed by the decoupling of the 6G network from end to end, providing corresponding capabilities and services. The management&orchestration plane realizes the flexible combination of functions across layers and planes, realizes the management and scheduling of multiple resources, and forms a complete service chain. This architecture focuses on multi-dimensional capability elements within the network. In [116], China Telecom proposed a ubiquitous and ultra-converge 6G network architecture, built upon the core concept of “three layers and three sectors”. The “three layers” consist of the infrastructure layer, the network function layer, and the network operating system layer, while the “three sectors” embody data integration, intelligent simplicity, and trustworthiness. This architecture leverages the intelligent simplicity sector to integrate the intelligent brain, connections, and facilities across the three layers, thereby enabling intelligent inclusiveness. China Unicom [117] envisioned a future 6G network as intelligent, converged, green, and trustworthy. Its architecture is structured vertically into the land-based, air-based, and space-based communication layers and horizontally into three domains: perception resource domain, function control domain, and service application domain. These layers and domains are interconnected by two chains: intelligent native and secure&trustworthy.
@@
Huawei [118] unveiled a task-centric AI architecture of 6G networks, asserting that 6G networks will introduce new resource dimensions and support the coordination of multi-dimensional resources. This design allows 6G networks to natively support AI and achieve a transformation from session-centric to task-centric operations. CICT [119, 120] put forward the concept of multi-network integration and ubiquitous intelligence, envisioning that 6G will realize global three-dimensional deep coverage, develop body area networks centered on human beings, form multi-layer coverage with both breadth and depth and learn the ubiquitous intelligence system with multi-network integration. OPPO [121] suggested that 6G would redefine human interaction and AI, by making AI a ubiquitous infrastructure. The 6G system architecture will emphasize the coordination across intelligence, performance, and flexibility, and integrate deeply AI capabilities into the user plane, control plane, and function plane. More authoritative and official, The IMT-2030 (6G) Promotion Group in China envisioned the 6G network architecture in [30], as an open and innovative platform for information services, offering capabilities that transcend mere connectivity. These capabilities encompass computing power networks, trusted security, sensing, data services, and AI-based network intelligent autonomy.
@@
Globally, Ericsson [122] has focused on cognitive networks by leveraging AI to achieve data-driven security and highly automated network operations. Qualcomm advocated that AI and ML would fundamentally transform the design and deployment of wireless communication and networking systems. As the leader of the EU’s 6G project Hexa-X, Nokia [123] stated that AI would be the primary driving force behind the technological transformation of the NR interface for 6G. The company is creating unique AI use cases and scenarios and developing foundational AI technologies. LG Electronics is focusing on the research and development of 6G AI technology, anticipating that 6G systems will usher in the era of the “Internet of Everything and Environment” powered by AI. Japan’s “Beyond 5G Promotion Strategy” project[124] aimed to introduce AI into network management, enabling network autonomy and driving network evolution towards intelligence.
@@
Academic research has also proposed various innovative ideas for shaping the future 6G network architecture. The authors of [125] put forth a 6G network design philosophy of “Human-Machine-Thing-Spirit”, emphasizing the inevitable trend of AI integration in 6G, leveraging cognitive enhancement and decision-making simulation to intelligently define network requirements, ensuring secure and reliable network transmission, and realizing intent-driven networks. In [126], a fusion of 6G RAN and AI was advocated, identifying four key characteristics: green, multi-dimensional, stereo, and full-scenario service. 6G would integrate communication, computation, control/caching, and AI to support full-application scenarios, and would be inherently secure[127]. The authors of [128] demonstrated a vision of integrating 6G and AI and studied implementation examples in wireless communications based on two classic AI algorithms: DL and RL. It was suggested in [33] that AI would arguably become the cornerstone of 6G with “Intelligent inclusion” serving as an essential feature of 6G, enabling the network to provide intelligent services to all types of end users. The authors of [129] promoted that AI would be used in wireless communications and change the top-level architecture of wireless networks. The authors of [130] emphasized that the NET4AI paradigm should leverage the capabilities of 6G networks, such as precise wireless sensing, high data-rate transmission, and ubiquitous connectivity, for intelligence distillation at the network edge, hereby promoting edge AI and enabling ubiquitous and trustworthy AI applications in wireless networks. Among them, federated edge learning is expected to be an important learning architecture over 6G. 6G networks and AI will be highly integrated, and customized AI services will be provided on the wireless network side[131]. 6G would enhance its progress in intelligence, especially in edge intelligence[132]. Among them, integrating DT and edge networks is a promising attempt. The 6G Flagship research program was initiated in [133], advocating for realizing context-aware intelligent services and applications for human and non-human users through AI.
@@
The industry has reached a core understanding of the intelligent design of 6G network architectures, emphasizing “distributed structure”, “endogenous intelligence”, and “integrated simplicity”. Future networks will evolve into advanced, integrated platforms that surpass mere communication connectivity services, offering multi-dimensional services and corresponding network capabilities beyond communication. They will cater to diverse businesses by offering end-to-end lifecycle management services throughout their entire process. Additionally, the network’s AI capability can be opened to the outside world, while external AI capability can be introduced into the network, enabling the crowdsourcing of AI capability.
@@
We elaborate on the NET4AI architecture, i.e., the support provided by the other five capabilities/services in the network (i.e., communication connectivity, data, computation, security, and management & orchestration) for the AI capabilities/services, with an emphasis on their interaction with AI, as illustrated in Figure 12. NET4AI will provide architectural support for the implementation of AI-native 6G, achieving a deep integration of the network and AI.